OpenStudy (anonymous):

8 combination 4 multiplied by 6 combination 1

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

\[8c4*6c1=\frac{ 8*7*6*5 }{ 4*3*2*1 }*\frac{ 6 }{ 1 }=?\]

OpenStudy (anonymous):

Thank you